# ToonifyGB: StyleGAN-based Gaussian Blendshapes for 3D Stylized Head Avatars

### v2

@@ Abstract
Introduction
With the advancement of 3D head reconstruction technologies, individuals are now able to personalize unique avatars for telepresence and virtual/augmented reality applications, which serve as an essential foundation for the rise of the metaverse. Considering user preferences and privacy concerns, the creation of stylized avatars is an important topic that deserves further research. Toonify (Pinkney and Adler 2020) is a StyleGAN-based method designed for 2D facial image stylization. In contrast to photo-realistic 3D head avatars, stylized 3D head avatars emphasize the expression of personal identity and the faithful transfer of target styles.
Blendshape is an efficient facial animation representation that synthesizes continuous and high-quality expressions by blending a set of 3D meshes, each corresponding to a specific facial expression. These facial shapes are synthesized by linearly blending the basis meshes using weighting coefficients. With the introduction of Neural Radiance Fields (NeRF) (Mildenhall et al. 2021), Gao et al. (Gao et al. 2022) and Zheng et al. (Zheng et al. 2022) incorporated the blendshape concept into NeRF, enabling avatar animation through the construction of a group of NeRF blendshapes that are linearly blended. Furthermore, the recently proposed 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) significantly improved rendering efficiency and delivered high-quality head reconstruction, outperforming NeRF in both speed and quality. Based on this, 3D Gaussian Blendshapes (3DGB) (Ma et al. 2024) successfully integrated the blendshape with Gaussian splatting, achieving real-time rendering and state-of-the-art performance in head reconstruction.
In contrast to previous works focused on photo-realistic 3D head avatar reconstruction, we propose ToonifyGB, a two-stage framework for synthesizing and animating 3D stylized head avatars. Given monocular video frames, Stage 1 adopt an improved StyleGAN to generate a more stable and less jittery stylized video, without requiring on fixed resolution and pre-aligned face cropping. In Stage 2, we begin with 3DGB to learn a neutral head model and a set of expression blendshapes, each represented as 3D Gaussians. Finally, by incorporating the facial tracker (Zielonka, Bolkart, and Thies 2022), we use the tracked motion parameters to drive ToonifyGB for animating 3D stylized head avatars.
In summary, the contributions of this work are as follows:
  • We propose ToonifyGB, an efficient two-stage framework that synthesizes 3D stylized head avatars from monocular videos using Gaussian blendshapes, supporting diverse styles and enabling real-time animation.
@@ Introduction
  • To the best of our knowledge, this work is the first to synthesize 3D stylized head avatars based on Gaussian blendshapes.
@@
Related Work
StyleGAN and Toonify
@@
StyleGAN (Karras, Laine, and Aila 2019; Karras et al. 2020) has been widely used to generate realistic facial images across diverse styles. Inversion of StyleGAN enables projecting real facial images into its latent space, allowing for subsequent edits such as adding glasses or changing hairstyle or age (Abdal, Qin, and Wonka 2019; Patashnik et al. 2021). To enhance inversion efficiency, methods such as pSp (Richardson et al. 2021) and e4e (Tov et al. 2021) employ encoders to directly project target faces into their corresponding latent codes. However, these methods often struggle to reconstruct fine image details, resulting in unsatisfactory reconstruction quality. To address these limitations, ReStyle (Alaluf, Patashnik, and Cohen-Or 2021) and HFGI (Wang et al. 2022) improve reconstruction fidelity by respectively predicting latent code residuals and correcting middle layer features. Nevertheless, these methods remain limited to aligned and cropped facial images to achieve effective editing and reconstruction.
@@
Recently, researchers (Pinkney and Adler 2020; Ojha et al. 2021; Jang et al. 2021; Yang et al. 2022a; Gal et al. 2022) have explored using StyleGAN for target‑domain image generation through transfer learning. Among these works, Toonify (Pinkney and Adler 2020) fine‑tunes the trained generator to blend realistic textures with toonified facial structures. In addition to image editing, StyleGAN has also been widely applied to video editing. Related studies have focused on enhancing video editing performance by employing temporal correlations in low‑dimensional latent codes (Fox et al. 2021), disentangling identity from facial attributes (Yao et al. 2021), incorporating sketch‑based branches (Liu et al. 2022), and tuning the generator to maintain temporal consistency (Tzaban et al. 2022). However, these methods typically rely on face alignment and cropping as preprocessing. Although StyleGAN3 (Karras et al. 2021) was introduced to support unaligned face inputs, a subsequent study (Alaluf et al. 2022) has shown that it struggles to encode facial features effectively without preprocessing, often resulting in structural artifacts. To overcome these limitations, methods such as VToonify (Yang et al. 2022b) and StyleGANEX (Yang et al. 2023) have been proposed to directly process videos beyond pre-aligned cropping. Nevertheless, these methods remain limited to 2D representations and have yet to be extended to 3D applications.
3D Head Avatar
@@
Since the introduction of NeRF (Mildenhall et al. 2021), implicit representation-based methods (Yenamandra et al. 2021; Zheng et al. 2022; Hong et al. 2022; Chan et al. 2022; Xu et al. 2023) for head reconstruction have achieved remarkable progress. 3DGS (Kerbl et al. 2023) has obtained a significant breakthrough in 3D reconstruction, further advancing the development of downstream applications such as 3D head modeling. Although several Gaussian-based methods (Qian et al. 2024; Xu et al. 2024; Chen et al. 2024; Ma et al. 2024; Xiang et al. 2024; Abdal et al. 2024; Kirschstein et al. 2024) have demonstrated high-quality head reconstruction and impressive rendering performance, they typically focus on photo-realistic avatars, with relatively limited exploration of avatar stylization. Stylized head avatars, characterized by geometric abstraction and artistic expression, differ significantly from the photo-realistic avatars synthesized by the aforementioned methods.
@@
Pre-trained 3D GANs (Wu et al. 2016) enable high-quality generation, making 3D head stylization possible. Although fine-tuning 3D generators for geometric and texture-based stylization has proven effective (Or-El et al. 2022; Jin et al. 2022; Abdal et al. 2023; Lan et al. 2023; Wang et al. 2023; Zhang et al. 2024), performing independent fine-tuning for each new style remains costly. Toonify3D (Jang et al. 2024) addressed this limitation by predicting facial surface normals using the proposed StyleNormal, enabling direct face stylization without additional fine-tuning. Similarly, DeformToon3D (Zhang et al. 2023) introduced StyleField to predict conditional 3D deformations, aligning NeRF representations in real space with style space to achieve geometric stylization and obviate per-style fine‑tuning. However, Toonify3D suffers from limited data diversity, and DeformToon3D cannot support novel-view animations, which limits their application scenarios.
Method
ToonifyGB Framework
@@
Given a monocular video input, ToonifyGB applies frame-by-frame stylization to generate the corresponding stylized frames. For inputs such as live streams or selfie videos, the face often occupies only a small portion of each frame, while the rest includes the hairstyle and upper body. Traditional methods (Karras, Laine, and Aila 2019; Karras et al. 2020) typically require face alignment, cropping, and editing before synthesizing the results back into the original frame. This process often introduces visual discontinuities at the seams, resulting in noticeable jitter in the output video. To address this issue, we adopt an improved StyleGAN model based on StyleGANEX (Yang et al. 2023) in Stage 1, enabling stable stylized video generation at the original resolution, as shown in Figure 1.
@@
To prepare the training data for Stage 2, we follow the method in (Zielonka, Bolkart, and Thies 2023; Ma et al. 2024), using the facial tracker from (Zielonka, Bolkart, and Thies 2022) to compute FLAME (Li et al. 2017) meshes, including a neutral head model and a set of expression blendshapes. This process also provides camera parameters $C$, joint and pose parameters $\Theta$, and expression coefficients $\{\psi_{k}\}$ for each frame. In addition to enabling facial expressions control, the FLAME model based on Principal Component Analysis (PCA) provides joint and pose parameters for controlling head, eyeball, eyelid, and jaw movements. As shown in Figure 1, we apply Linear Blend Skinning (LBS) to transform the Gaussian model based on the extracted joint and pose parameters. The transformation is defined as:
@@
The transformed Gaussian model is then rendered in real-time as a 3D stylized head avatar using Gaussian Splatting. Finally, by integrating the camera parameters $C$, we enable novel-view rendering and animation.
Figure 2: Visualization of stylized video generation results on the INSTA (Zielonka, Bolkart, and Thies 2023) and NeRFBlendShape (Gao et al. 2022) datasets.
Stylized Video Generation
As shown in StyleGANEX (Yang et al. 2023), manipulating feature maps at different layers of StyleGAN leads to different spatial effects in the generated faces. Specifically, while shifting or rotating the feature maps in deeper layers (i.e., Layer 7) produces consistent global transformations, similar operations in shallow layers (i.e., Layer 1) fail to preserve facial structure due to the low spatial resolution of the $4\times 4$ feature map, causing blurring and loss of detail. To address this limitation, we adopt StyleGANEX (Yang et al. 2023), an enhanced variant of StyleGAN2 (Karras, Laine, and Aila 2019), which increases the spatial resolution of shallow feature maps (Layers 1–7) to $32\times 32$. This improvement enables finer control over facial geometry and enhances the generation quality for unaligned faces.
@@
Gaussian Blendshapes Synthesis
We represent all Gaussian head avatars using 3D Gaussians. As described in 3DGS (Kerbl et al. 2023), each Gaussian has some basic properties including Gaussian center $\mu$, scale $s$, color $c$, opacity $\alpha$, and rotation $q$. Based on 3DGB (Ma et al. 2024), our Gaussian blendshape representation consists of a neutral model $B_{0}$ and a set of $n$ expression blendshapes ${B_{1},B_{2},\ldots,B_{n}}$. Each Gaussian in the neutral model $B_{0}$ has a set of blend weights $w$ to control joint and pose. In addition, each Gaussian in an expression blendshape $B_{k}$ corresponds one‑to‑one to a Gaussian in the neutral model $B_{0}$. The difference between $B_{k}$ and $B_{0}$ is defined as the difference in their corresponding Gaussian properties: $\Delta B_{k}=B_{k}-B_{0}$. The expression of Gaussian head avatar $B^{\psi}$ can be computed as follows:
@@
Since the FLAME meshes and blendshape models do not include interior mouth components such as teeth, we adopt the method of 3DGB (Ma et al. 2024) by defining a separate set of Gaussians for the mouth $B_{m}$. The properties of these mouth Gaussians are not affected by expression changes, they only move with the jaw joint in the FLAME model. The mouth Gaussians for the head avatar, $B_{m}^{*}$, are computed via linear blend skinning (LBS) as:
@@
Loss Function
We adopt the loss function from 3DGB (Ma et al. 2024), and define the total loss as follows:
@@
$L_{\alpha}=\frac{1}{F}\sum_{i=1}^{F}\left\|I_{\alpha}^{i}-M_{h}^{i}\right\|_{2}.$ (6)
@@
Experiments
Baseline and Dataset
Due to the current lack of methods for synthesizing 3D stylized head avatars using Gaussian blendshapes, we compare our method against the following state-of-the-art methods for photo-realistic 3D head avatar synthesis: INSTA (Zielonka, Bolkart, and Thies 2023), PointAvatar (Zheng et al. 2023), FLARE (Bharadwaj et al. 2023), SplattingAvatar (Shao et al. 2024), FlashAvatar (Xiang et al. 2024), and 3DGB (Ma et al. 2024). Notably, 3DGB shares a similar architecture with ours but focuses on photo-realistic avatar synthesis and does not support the synthesis of diverse stylized avatars.
We evaluate both our method and state-of-the-art photo-realistic avatar synthesis methods using six videos from the INSTA (Zielonka, Bolkart, and Thies 2023) dataset. Each video is cropped and resized to $512\times 512$ resolution, with sequence lengths ranging from 1,000 to 4,000 frames. Following the method of 3DGB (Ma et al. 2024), we retain the final 350 frames of each video for testing. Both 3DGB (Ma et al. 2024) and our method apply the same preprocessing pipeline (Zielonka, Bolkart, and Thies 2022, 2023), including background removal and FLAME parameter extraction.
Evaluation Metrics
We employ two metrics to evaluate video stabilization: Inter-frame Transformation Fidelity (ITF) (Morimoto and Chellappa 1998; Marcenaro, Vernazza, and Regazzoni 2001; Xu et al. 2012) and Inter-frame Similarity Index (ISI) (Guilluy, Beghdadi, and Oudre 2018; James, Jain, and Rajwade 2023). ITF measures the inter-frame Peak Signal-to-Noise Ratio (PSNR) in dB based on the mean squared error. The intuitive idea of ITF is that a more stable video (i.e., less jittery) will have greater similarity between adjacent frames compared to an unstable version of the same video. ISI computes the average Structural Similarity (SSIM) between adjacent frames across the video. Higher ISI values indicate greater perceptual similarity between frames, leading to improved visual comfort for viewers.
For 3D head avatar synthesis, we evaluate the performance of our method and state-of-the-art methods for photo-realistic avatar synthesis using standard evaluation metrics (Zhang et al. 2018; Ma et al. 2024), including Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM). In addition, we record the training time (in minutes) and the rendering speed (in frames per second, fps) for each method. In the ablation study, we additionally adopt the Learned Perceptual Image Patch Similarity (LPIPS) metric to better capture perceptual differences between the synthesized avatars and the ground truth.
Figure 3: Visualization of stylized video generation: We present details of the real head from the input video, and the “Arcane” stylized head generated by our method. From left to right, the results for the data samples “bala” and “wojtek_1” are shown.
Implementation Details
To ensure a fair performance comparison, the training and testing of all methods are performed on a single RTX 4090 GPU. Our methods are implemented in Python using the PyTorch framework. For 2D stylized video generation, we use the pre-trained models provided by StyleGANEX (Yang et al. 2023). For training the 3D stylized head avatars, we employ the Adam optimizer (Kingma and Ba 2015), setting the initial learning rates of the Gaussian properties $\{\mathbf{x}_{k},\alpha_{k},\mathbf{s}_{k},\mathbf{q}_{k},SH_{k}\}$ to $3.2\times 10^{-7}$, $5\times 10^{-5}$, $5\times 10^{-4}$, $1\times 10^{-4}$, and $1.25\times 10^{-3}$, respectively. Following 3DGB (Ma et al. 2024), the initial number of sampled Gaussians is 50k for the neutral head model and 14k for the mouth interior.
Table 1: Duration (in seconds) of the input videos, and inference time (in seconds) of our method.
Dataset justin malte_1 nf_01 bala wojtek_1 person_0004

### v3

@@ Abstract
1 Introduction
With the advancement of 3D head reconstruction technologies, individuals can now personalize unique avatars for telepresence and virtual/augmented reality applications, which serve as a crucial foundation for the rise of the metaverse. Considering user preferences and privacy concerns, the creation of stylized avatars has become an important research topic. Toonify [36], a StyleGAN-based method, was designed for 2D facial image stylization, presenting the potential of translating real portraits into stylized 2D images. While such methods focus on 2D images, recent advances in 3D head reconstruction have mainly targeted photo-realistic avatars. In contrast, stylized 3D head avatars emphasize personal identity and the faithful transfer of artistic styles.
Blendshapes are an efficient facial animation representation that synthesize continuous and high-quality expressions by blending a set of 3D meshes, each corresponding to a specific facial expression. These facial shapes are synthesized by linearly blending the basis meshes using weighting coefficients. With the introduction of Neural Radiance Fields (NeRF) [31], Gao et al. [11] and Zheng et al. [57] incorporated the blendshape concept into NeRF, enabling avatar animation through a group of NeRF blendshapes that are linearly blended. Furthermore, the recently proposed 3D Gaussian Splatting (3DGS) [22] significantly improved rendering efficiency and achieved higher-quality head reconstruction, outperforming NeRF in both speed and quality. Building on this, 3D Gaussian Blendshapes (3DGB) [29] successfully integrated blendshapes with Gaussian splatting, achieving real-time rendering and state-of-the-art performance in head reconstruction.
In contrast to previous works focused on photo-realistic 3D head avatar reconstruction, we propose ToonifyGB, a two-stage framework for synthesizing and animating 3D stylized head avatars. Given monocular video frames, Stage 1 adopts an improved StyleGAN to generate a more stable and less jittery stylized video, without requiring fixed resolution or pre-aligned face cropping. In Stage 2, we build upon 3DGB to learn a neutral head model and a set of expression blendshapes, each represented as 3D Gaussians. Finally, by incorporating a facial tracker [59], ToonifyGB uses the tracked motion parameters to animate 3D stylized head avatars.
The contributions of this work are as follows:
  • We propose ToonifyGB, an efficient two-stage framework that synthesizes 3D stylized head avatars from monocular videos using Gaussian blendshapes, supporting diverse styles with real-time animation.
@@ 1 Introduction
  • To the best of our knowledge, this work is the first to synthesize 3D stylized head avatars using Gaussian blendshapes.
@@
2 Related Work
2.1 StyleGAN and Toonify
@@
StyleGAN [19, 20] has been widely used to generate realistic facial images across diverse styles. Inversion of StyleGAN enables projecting real facial images into its latent space, allowing subsequent edits such as adding glasses or changing hairstyles or age [2, 35]. To enhance inversion efficiency, methods such as pSp [38] and e4e [40] employ encoders to directly project target faces into their corresponding latent codes. However, these methods often struggle to reconstruct fine image details, resulting in unsatisfactory reconstruction quality. To address these limitations, ReStyle [4] and HFGI [43] improve reconstruction fidelity by respectively predicting latent code residuals and correcting intermediate features. Nevertheless, these methods remain limited to aligned and cropped facial images for effective editing and reconstruction.
@@
Recently, researchers [36, 33, 15, 49, 10] have explored the use of StyleGAN for target-domain image generation through transfer learning. Among these works, Toonify [36] fine‑tunes the trained generator to blend realistic textures with toonified facial structures. In addition to image editing, StyleGAN has also been widely applied to video editing. Related studies have focused on enhancing video editing performance by employing temporal correlations in low‑dimensional latent codes [9], disentangling identity from facial attributes [52], incorporating sketch‑based branches [28], and tuning the generator to maintain temporal consistency [41]. However, these methods typically rely on face alignment and cropping as preprocessing. Although StyleGAN3 [18] was introduced to support unaligned face inputs, a subsequent study [5] has shown that it struggles to encode facial features effectively without preprocessing, often resulting in structural artifacts. To overcome these limitations, methods such as VToonify [50] and StyleGANEX [51] have been proposed to directly process videos beyond pre-aligned cropping. Nevertheless, these methods remain limited to 2D representations and have yet to be extended to 3D applications.
2.2 3D Head Avatar
@@
Since the introduction of NeRF [31], implicit representation-based methods [53, 57, 13, 7, 48] for head reconstruction have achieved remarkable progress. 3DGS [22] has obtained a significant breakthrough in 3D reconstruction, further advancing the development of downstream applications such as 3D head modeling. Although several Gaussian-based methods [37, 47, 8, 29, 45, 3, 24] have demonstrated high-quality head reconstruction and impressive rendering performance, they typically focus on photo-realistic avatars, with relatively limited exploration of avatar stylization. Stylized head avatars, characterized by geometric abstraction and artistic expression, differ significantly from the photo-realistic avatars synthesized by the aforementioned methods.
@@
Pre-trained 3D GANs [44] enable high-quality generation, making 3D head stylization possible. Although fine-tuning 3D generators for geometric and texture-based stylization has proven effective [34, 17, 1, 25, 42, 54], performing independent fine-tuning for each new style remains costly. Toonify3D [16] addressed this limitation by predicting facial surface normals using the proposed StyleNormal, enabling direct face stylization without additional fine-tuning. Similarly, DeformToon3D [55] introduced StyleField to predict conditional 3D deformations, aligning NeRF representations in real space with style space to achieve geometric stylization and obviate per-style fine‑tuning. However, Toonify3D suffers from limited data diversity, and DeformToon3D cannot support novel-view animations, which limits their application scenarios.
3 Method
3.1 ToonifyGB Framework
@@
Given a monocular video input, ToonifyGB applies frame-by-frame stylization to generate the corresponding stylized frames. For inputs such as live streams or selfie videos, the face often occupies only a small portion of each frame, while the rest includes the hairstyle and upper body. Traditional methods [19, 20] typically require face alignment, cropping, and editing before synthesizing the results back into the original frame. This process often introduces visual discontinuities at the seams, resulting in noticeable jitter in the output video. To address this issue, we adopt an improved StyleGAN model based on StyleGANEX [51] in Stage 1, enabling stable stylized video generation at the original resolution, as shown in Figure 2.
@@
To prepare the training data for Stage 2, we follow the method in [60, 29], using the facial tracker from [59] to compute FLAME [26] meshes, including a neutral head model and a set of expression blendshapes. This process also provides camera parameters $C$, joint and pose parameters $\Theta$, and expression coefficients $\{\psi_{k}\}$ for each frame. In addition to enabling facial expressions control, the FLAME model based on Principal Component Analysis (PCA) provides joint and pose parameters for controlling head, eyeball, eyelid, and jaw movements. As shown in Figure 2, we apply Linear Blend Skinning (LBS) to transform the Gaussian model based on the extracted joint and pose parameters. The transformation is defined as:
@@
The transformed Gaussian model is then rendered in real-time as a 3D stylized head avatar using Gaussian Splatting. Finally, by integrating the camera parameters, we enable novel-view rendering and animation.
Figure 3: Visualization of stylized video generation results in “Arcane” and “Pixar” styles on the INSTA [60] and NeRFBlendShape [11] datasets, covering both male and female subjects.
3.2 Stylized Video Generation
As shown in StyleGANEX [51], manipulating feature maps at different layers of StyleGAN leads to different spatial effects in the generated faces. Specifically, while shifting or rotating the feature maps in deeper layers (i.e., Layer 7) produces consistent global transformations, similar operations in shallow layers (i.e., Layer 1) fail to preserve facial structure due to the low spatial resolution of the $4\times 4$ feature map, causing blurring and loss of detail. To address this limitation, we adopt StyleGANEX [51], an enhanced variant of StyleGAN2 [19], which increases the spatial resolution of shallow feature maps (Layers 1–7) to $32\times 32$. This improvement enables finer control over facial geometry and enhances the generation quality for unaligned faces.
@@ 3.2 Stylized Video Generation
3.3 Gaussian Blendshapes Synthesis
We represent all Gaussian head avatars using 3D Gaussians. As described in 3DGS [22], each Gaussian has some basic properties including Gaussian center $\mu$, scale $s$, color $c$, opacity $\alpha$, and rotation $q$. Based on 3DGB [29], our Gaussian blendshape representation consists of a neutral model $B_{0}$ and a set of $n$ expression blendshapes $B_{1},B_{2},\ldots,B_{n}$}. Each Gaussian in the neutral model $B_{0}$ has a set of blend weights $w$ to control joint and pose. In addition, each Gaussian in an expression blendshape $B_{k}$ corresponds one‑to‑one to a Gaussian in the neutral model $B_{0}$. The difference between $B_{k}$ and $B_{0}$ is defined as the difference in their corresponding Gaussian properties: $\Delta B_{k}=B_{k}-B_{0}$. The expression of Gaussian head avatar $B^{\psi}$ can be computed as follows:
@@ 3.3 Gaussian Blendshapes Synthesis
Since the FLAME meshes and blendshape models do not include interior mouth components such as teeth, we adopt the method of 3DGB [29] by defining a separate set of Gaussians for the mouth $B_{m}$. The properties of these mouth Gaussians are not affected by expression changes, they only move with the jaw joint in the FLAME model. The mouth Gaussians for the head avatar, $B_{m}^{*}$, are computed via linear blend skinning (LBS) as:
@@
Figure 4: Visualization of stylized video generation results: We present details of the real head from the input video, and the “Arcane” stylized head generated by our method. From left to right, the results for the video samples “bala” and “wojtek_1” are shown.
3.4 Loss Function
We adopt the loss function from 3DGB [29], and define the total loss as follows:
@@ 3.4 Loss Function
$L_{\alpha}=\frac{1}{F}\sum_{i=1}^{F}\frac{1}{|P|}\sum_{p\in P}\big(I_{\alpha}^{i}(p)-M_{h}^{i}(p)\big)^{2}.$ (6)
@@
Table 1: Video durations and inference times: Duration (in seconds) of the input videos, and inference time (in seconds) of our method.
Video samples justin malte_1 nf_01 bala wojtek_1 person_0004
@@
4 Experiments
4.1 Baselines
Due to the current lack of methods for synthesizing 3D stylized head avatars using Gaussian blendshapes, we compare our method against the following state-of-the-art methods for photo-realistic 3D head avatar synthesis: INSTA [60], PointAvatar [58], FLARE [6], SplattingAvatar [39], FlashAvatar [45], and 3DGB [29]. Notably, 3DGB shares a similar architecture with ours but focuses on photo-realistic avatar synthesis and does not support the synthesis of diverse stylized avatars.